Calculate the perpetual equivalent annual cost (years 1 to ∞) of $850,000 now and $750,000, 4 years from now at an interest rate of 12% per year.

The perpetual equivalent annual cost is determined to be $
Answer #1

Solution

The perpetual equivalent annual cost of $850,000 now and $750,000 4 years from now at an interest rate of 12% per annum is calculated as below,

P= [$850,000 + $750,000 (P/F, 12%, 4)] (A/P,12%,∞)

= [850,000 + 750,000 (0.6355)] x12%

= (850,000 + 476625) x0.12=$159195

Note – (A/P, i, ∞) = i

Hence, the perpetual equivalent annual cost of $850,000 now and $750,000 4 years from now at an interest rate of 12% per annum is $159195